# Atomic sheaves on hyper-Kähler manifolds via Bridgeland moduli spaces

### v2

@@ 5.4. Twisted atomic bundles
Given a hyper-Kähler manifold $X$ and a Brauer class $\alpha\in\mathrm{H}^{2}(X,\operatorname{\mathcal{O}}_{X}^{*})$, we consider the twisted derived category $\mathrm{D}^{\mathrm{b}}(X,\alpha)$. In [12, Section 4.4], the author extends the theory of atomicity to the case of twisted objects in $\mathrm{D}^{\mathrm{b}}(X,\alpha)$ and proves some properties of twisted atomic sheaves in a specific setting.
On the other hand, for a twisted vector bundle on $X$, the notion of projectively hyperholomorphic bundles is introduced in [75], which has well-behaved deformation properties. We refer to [53, Section 6] for more details.
@@
We fix a pair of coprime integers $a,b$ and set $n:=a^{2}+b^{2}+1$. Let $X$ be a general GM fourfold and $Y\subset X$ be a general hyperplane section. By Theorem 5.8, we have an immersed atomic Lagrangian submanifold $r\colon M^{Y}_{\sigma_{Y}}(a,b)\to M^{X}_{\sigma_{X}}(a,b)$. Up to the $\widetilde{\mathrm{GL}}^{+}_{2}(\mathbb{R})$-action, we may assume that $\sigma_{X}$ and $\sigma_{Y}$ are rational with the same central charge. By Theorem 3.7 and Lemma 3.8, we can choose a non-rigid connected component $L\subset M^{Y}_{\sigma_{Y}}(a,b)$ such that $r$ maps $L$ birationally to its image $Z$. Then, by Lemma 4.9, the restriction of $r$ on $L$, denoted also by $r$, is a connected immersed atomic Lagrangian submanifold.
@@
Recall that $M^{X}_{\sigma_{X}}(a,b)$ forms an open dense subset of the moduli space of projective hyper-Kähler manifolds of degree $22n-$ 2 and divisibility $n-1$ when $X$ varies in an open dense subset of the moduli space of GM fourfolds (cf. [66, Theorem 1.7]). By the Torelli Theorem [32, Theorem 8.1], we can take $X$ such that $\mathrm{H}^{1,1}_{\mathbb{Z}}(M^{X}_{\sigma_{X}}(a,b))=\langle h,f\rangle$, where $h$ is the natural polarization and $f$ induces a Lagrangian fibration $M^{X}_{\sigma_{X}}(a,b)\to\mathbb{P}^{n}$ such that $q(h,f)=d$ for a positive integer $d$. When $d\gg 0$, the movable cone of $M^{X}_{\sigma_{X}}(a,b)$ equals to the nef cone of $M^{X}_{\sigma_{X}}(a,b)$ by [55, Corollary 2.7] and [61, Lemma 4.3].
@@
Assume that $d$ is sufficiently large and coprime with $n-1=a^{2}+b^{2}$. Then, according to [62, Proposition 5.12] or [52, Theorem 7.13], we can take $X$ such that the fibration $M^{X}_{\sigma_{X}}(a,b)\to\mathbb{P}^{n}$ is isomorphic to a Tate–Shafarevich twist of $M_{S}(0,\mathcal{L},1-n)\to|\mathcal{L}|$, where $S$ is a K3 surface with $\operatorname{Pic}(S)=\mathbb{Z}\mathcal{L}$. Then, using [12, Theorem 6.2.5], there is a Brauer class $\alpha\in\mathrm{H}^{2}(M^{X}_{\sigma_{X}}(a,b),\operatorname{\mathcal{O}}^{*}_{M^{X}_{\sigma_{X}}(a,b)})$ and a twisted Poincaré sheaf on $M^{X}_{\sigma_{X}}(a,b)\times_{\mathbb{P}^{n}}M_{S}(0,\mathcal{L},1-n)$ such that the corresponding Fourier–Mukai functor $\Phi$ is an equivalence
@@
In particular, for any $\operatorname{\mathcal{F}}\in\operatorname{Pic}^{0}(L)$, its image $\Phi(r_{*}\operatorname{\mathcal{F}})\in\mathrm{D}^{\mathrm{b}}(M_{S}(0,\mathcal{L},1-n),\alpha)$ is a twisted atomic object by [12, Proposition 4.4.7].
Now, we claim $f^{n}.r_{*}[L]\neq 0$, which implies that $Z\to\mathbb{P}^{n}$ is finite. Indeed, by Corollary 5.2, we have $h^{\perp}=\ker(r^{*})$. As $q(h,f)=d\neq 0$, we see $f\notin\ker(r^{*})$. By the atomicity, $r^{*}h$ spans $\operatorname{im}(r^{*})$. Then $r^{*}f=t\cdot r^{*}h\in\mathrm{H}^{2}(L,\mathbb{Q})$ for a non-zero $t\in\mathbb{Q}$. Hence, we have $f^{n}.r_{*}[L]=t^{n}\cdot h^{n}.r_{*}[L]\neq 0$.
Since $r$ is finite, $r_{*}\operatorname{\mathcal{F}}$ is a rank one Cohen–Macaulay sheaf on $Z$ for any $\operatorname{\mathcal{F}}\in\operatorname{Pic}^{0}(L)$ by [40, Corollary 2.73]. Then, from the fact that $Z\to\mathbb{P}^{n}$ is finite, $\Phi(r_{*}\operatorname{\mathcal{F}})$ is a twisted atomic bundle by [12, Proposition 5.2.6, Proposition 6.5.4]. As $Z$ is integral and $r_{*}\operatorname{\mathcal{F}}$ is Cohen–Macaulay of rank one on $Z$, the proof of [12, Theorem 6.5.6] applies in our case. Then we know that $\Phi(r_{*}\operatorname{\mathcal{F}})$ is stable with respect to any suitable polarization.
@@
Remark 5.12.
We expect that the non-rigid bundle constructed in Theorem 1.2 has a $20$-dimensional deformation space. This can be deduced from the connectedness of $M^{Y}_{\sigma_{Y}}(a,b)$ and the fact that $b_{1}(M^{Y}_{\sigma_{Y}}(a,b))=b_{3}(Y)=20$. We anticipate that the arguments used in [47] for cubic threefolds can be applied in this case as well.
@@ 6.1. Double dual EPW sextics
Recall that by [17, Theorem 3.10], each smooth GM variety $X$ of dimension $\dim X\geq 3$ is uniquely determined by its Lagrangian data
@@
as in [56, Section 2]. See also [17, Appendix B] for an overview. We set $\mathsf{Y}_{A(X)}:=\mathsf{Y}_{A(X)}^{\geq 1}$ and $\mathsf{Y}_{A(X)^{\perp}}:=\mathsf{Y}_{A(X)^{\perp}}^{\geq 1}$, which are called the Eisenbud–Popescu–Walter (EPW) sextic and the dual EPW sextic associated with $X$, respectively.
Moreover, by [58, Section 1.2] and [19, Theorem 5.2], as $A(X)$ does not contain decomposable vectors [^6] This means $\mathbb{P}(A(X))\cap\operatorname{Gr}(3,V_{6})=\varnothing$, which holds for $A(X)$ when $X$ is smooth and $\dim X\geq 3$ by [17, Theorem 3.16]., we have natural double coverings
@@
Since $X$ and $Y$ are both ordinary, by [17, Propoition 3.14], the Lagrangian data of $X$ and $Y$ are $(V_{6},V_{5},A(X))$ and $(V_{6},V_{5},A(Y))$ respectively, with $\dim(A(X)\cap A(Y))=9$. Therefore, part (1) follows directly from the above definition of EPW varieties.
@@
Consequently, $\iota$ is finite and unramified in codimension one. Thus, the natural map $W\to\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ is also unramified in codimension one. Then, by the smoothness of $\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ and [29, Theorem 1.6], it follows that $W\to\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ is unramified, and so is $\iota$.
@@
Moreover, since $W\to\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ is unramified and finite of degree two, by [71, Tag 0B3D] and [71, Tag 00R4], it is flat and hence an étale double covering. In particular, $W$ is smooth. Then, by the uniqueness part of [19, Theorem 5.2], we have $W\cong\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}$ and (2) follows. ∎
@@
It is suggested in [36, Section 5.1] that $i$ is injective, implying that $i$ is an embedding. Nevertheless, this is equivalent to $\varnothing=\mathsf{Y}^{\geq 2}_{A(Y)^{\perp}}\cap\mathsf{Y}_{A(X)^{\perp}}^{\geq 2}\subset\mathsf{Y}_{A(X)^{\perp}}$, which is impossible as explained in [24, Lemma 3.2]. See Remark 6.15 for an alternative proof of the non-injectivity of $i$.

### v3

@@ 5.4. Twisted atomic bundles
Given a hyper-Kähler manifold $X$ and a Brauer class $\alpha\in\mathrm{H}^{2}(X,\operatorname{\mathcal{O}}_{X}^{*})$, we consider the twisted derived category $\mathrm{D}^{\mathrm{b}}(X,\alpha)$. In [12, Section 4.4] and [13, Section 2], the author extends the theory of atomicity to the case of twisted objects in $\mathrm{D}^{\mathrm{b}}(X,\alpha)$ and proves some properties of twisted atomic sheaves in a specific setting.
On the other hand, for a twisted vector bundle on $X$, the notion of projectively hyperholomorphic bundles is introduced in [80], which has well-behaved deformation properties. We refer to [58, Section 6] for more details.
@@
We fix a pair of coprime integers $a,b$ and set $n:=a^{2}+b^{2}+1$. Let $X$ be a general GM fourfold and $j\colon Y\hookrightarrow X$ be a general hyperplane section. By Theorem 5.8, we have an immersed atomic Lagrangian submanifold $r\colon M^{Y}_{\sigma_{Y}}(a,b)\to M^{X}_{\sigma_{X}}(a,b)$. Up to the $\widetilde{\mathrm{GL}}^{+}_{2}(\mathbb{R})$-action, we may assume that $\sigma_{X}$ and $\sigma_{Y}$ are rational with the same central charge. From Theorem 3.7 and Lemma 3.8, we can choose a connected component $L\subset M^{Y}_{\sigma_{Y}}(a,b)$ such that $r$ maps $L$ birationally to its image $Z$ and hence $r(L)=Z$ is not contained in the fixed locus of the natural involution on $M^{X}_{\sigma_{X}}(a,b)$. The restriction of $r$ on $L$, denoted also by $r$, is a connected immersed atomic Lagrangian submanifold by Lemma 4.9.
We claim that $\mathrm{H}^{1}(L,\operatorname{\mathcal{O}}_{L})\neq 0$, i.e. $L$ is non-rigid. Assume for the contradiction that $\mathrm{H}^{1}(L,\operatorname{\mathcal{O}}_{L})$ vanishes. Up to equivalences of Kuznetsov components, we may assume that $X$ is special and $Y$ is not the branch divisor of $X$ by [46, Theorem 1.6] and [45, Lemma 3.8]. Since $\mathrm{H}^{1}(L,\operatorname{\mathcal{O}}_{L})=0$, the Albanese variety of $L$ is trivial. In particular, the Abel–Jacobi map
$L\to\mathrm{J}(Y)$
defined by using the second Chern class is trivial, where $\mathrm{J}(Y)$ is the intermediate Jacobian of $Y$. Then [28, Theorem 4.1(ii)] implies that $c_{2}(E)=c_{2}(E^{\prime})\in\mathrm{CH}^{2}(Y)$ for any two points $[E],[E^{\prime}]\in L$. Using [28, Theorem 4.1(i)], this further implies
$\mathrm{ch}(E)=\mathrm{ch}(E^{\prime})\in\mathrm{CH}^{*}(Y)_{\mathbb{Q}}\quad\text{and}\quad\mathrm{ch}(\mathrm{pr}_{X}(j_{*}E))=\mathrm{ch}(\mathrm{pr}_{X}(j_{*}E^{\prime}))\in\mathrm{CH}^{*}(X)_{\mathbb{Q}}.$
Therefore, from [27, Proposition 6.6], we know that $r(L)\subset M^{X}_{\sigma_{X}}(a,b)$ is a Lagrangian constant cycle in the sense of [83], i.e. the set of rational equivalence classes of points of $r(L)$ in $\mathrm{CH}_{0}(M^{X}_{\sigma_{X}}(a,b))$ is a singleton. Now, we consider the diagram
${{L(a,b)}}$${{M^{X}_{\sigma_{X}}(a,b)}}$${V\subset|\operatorname{\mathcal{O}}_{X}(H)|}$$\scriptstyle{q}$$\scriptstyle{p}$
as in the proof of [27, Theorem 5.8], where $V\subset|\operatorname{\mathcal{O}}_{X}(H)|$ is the open subset parameterizing hyperplane sections of $X$ such that the moduli space of class $a\lambda_{1}+b\lambda_{2}$ is smooth and $p$ is the relative moduli space which is smooth projective. Moreover, $q|_{p^{-1}(s)}$ is the morphism constructed in Theorem 3.7(1) for each $s\in V$. We denote by $0\in V$ the point corresponding to $Y$. Since $X$ is general, there is also a point $1\in V$ corresponding to the branch divisor of $X$. Up to taking the Stein factorization, we can assume that $p$ has connected fibers. Moreover, by replacing $L(a,b)$ with the component containing $L\subset p^{-1}(0)$, we can assume that $L(a,b)$ is irreducible. Since $p$ is a smooth proper family, we have $\mathrm{H}^{1}(p^{-1}(s),\operatorname{\mathcal{O}}_{p^{-1}(s)})=\mathrm{H}^{1}(L,\operatorname{\mathcal{O}}_{L})=0$ for each $s\in V$. In particular, by the same argument as above, we see that $q(p^{-1}(s))$ is a Lagrangian constant cycle in $M^{X}_{\sigma_{X}}(a,b)$ for each $s\in V$. Therefore, applying [83, Theorem 1.3(i)], we conclude that $q(L(a,b))$ has dimension $n=\frac{1}{2}\dim M^{X}_{\sigma_{X}}(a,b)$. Since $L(a,b)$ is irreducible, $q(L(a,b))$ is also irreducible. Therefore, we have $q(L(a,b))=q(p^{-1}(s))$ for each $s\in V$ as $q(p^{-1}(s))\subset q(L(a,b))$ and they have the same dimension. However, this leads to a contradiction, since $q(p^{-1}(1))$ lies in the fixed locus of the natural involution on $M^{X}_{\sigma_{X}}(a,b)$, while $q(p^{-1}(0))=q(L)$ does not, by our choice of $L$. This proves the claim.
@@
Now, back to the proof of our theorem. Recall that $M^{X}_{\sigma_{X}}(a,b)$ forms an open dense subset of the moduli space of projective hyper-Kähler manifolds of degree $22n-$ 2 and divisibility $n-1$ when $X$ varies in an open dense subset of the moduli space of GM fourfolds (cf. [71, Theorem 1.7]). By the Torelli Theorem [35, Theorem 8.1], we can take $X$ such that $\mathrm{H}^{1,1}_{\mathbb{Z}}(M^{X}_{\sigma_{X}}(a,b))=\langle h,f\rangle$, where $h$ is the natural polarization and $f$ induces a Lagrangian fibration $M^{X}_{\sigma_{X}}(a,b)\to\mathbb{P}^{n}$ such that $q(h,f)=d$ for a positive integer $d$. When $d\gg 0$, the movable cone of $M^{X}_{\sigma_{X}}(a,b)$ equals to the nef cone of $M^{X}_{\sigma_{X}}(a,b)$ by [60, Corollary 2.7] and [66, Lemma 4.3].
@@
Assume that $d$ is sufficiently large and coprime with $n-1=a^{2}+b^{2}$. Then, according to [67, Proposition 5.12] or [57, Theorem 7.13], we can take $X$ such that the fibration $M^{X}_{\sigma_{X}}(a,b)\to\mathbb{P}^{n}$ is isomorphic to a Tate–Shafarevich twist of $M_{S}(0,\mathcal{L},1-n)\to|\mathcal{L}|$, where $S$ is a K3 surface with $\operatorname{Pic}(S)=\mathbb{Z}\mathcal{L}$. Then, using [13, Theorem 3.3], there is a Brauer class $\alpha\in\mathrm{H}^{2}(M^{X}_{\sigma_{X}}(a,b),\operatorname{\mathcal{O}}^{*}_{M^{X}_{\sigma_{X}}(a,b)})$ and a twisted Poincaré sheaf on $M^{X}_{\sigma_{X}}(a,b)\times_{\mathbb{P}^{n}}M_{S}(0,\mathcal{L},1-n)$ such that the corresponding Fourier–Mukai functor $\Phi$ is an equivalence
@@
In particular, for any $\operatorname{\mathcal{F}}\in\operatorname{Pic}^{0}(L)$, its image $\Phi(r_{*}\operatorname{\mathcal{F}})\in\mathrm{D}^{\mathrm{b}}(M_{S}(0,\mathcal{L},1-n),\alpha)$ is a twisted atomic object by [13, Corollary 3.10].
@@
As $r$ is finite, $r_{*}\operatorname{\mathcal{F}}$ is a rank one Cohen–Macaulay sheaf on $Z$ for any $\operatorname{\mathcal{F}}\in\operatorname{Pic}^{0}(L)$ by [44, Corollary 2.73]. Then, from the fact that $Z\to\mathbb{P}^{n}$ is finite, $\Phi(r_{*}\operatorname{\mathcal{F}})$ is a twisted atomic bundle by [13, Proposition 6.2]. So [13, Theorem 6.6] implies that $\Phi(r_{*}\operatorname{\mathcal{F}})$ is stable with respect to any suitable polarization.
[…]
We expect that the bundle constructed in Theorem 1.2 has a $20$-dimensional deformation space. This can be probably deduced from the connectedness of $M^{Y}_{\sigma_{Y}}(a,b)$ and the fact that $b_{1}(M^{Y}_{\sigma_{Y}}(a,b))=b_{3}(Y)=20$, together with Lemma 5.12. We anticipate that the arguments used in [51] for cubic threefolds can be applied in this case as well.
Lemma 5.12.
Let $M$ be a projective hyper-Kähler manifold and $r\colon L\to M$ be an immersed connected Lagrangian submanifold. Then for any line bundle $\mathcal{L}$ on $L$, we have
$\mathrm{H}^{1}(L,\operatorname{\mathcal{O}}_{L})\subset\operatorname{Ext}_{M}^{1}(r_{*}\mathcal{L},r_{*}\mathcal{L}).$
Moreover, if $r$ is a closed embedding away from a closed subset of codimension $\geq 2$ in $L$, then we have
$\dim\operatorname{Ext}_{M}^{1}(r_{*}\mathcal{L},r_{*}\mathcal{L})\leq\dim\mathrm{H}^{1}(L,\mathbb{C})$ (14)
and the equality holds if there exists a smooth projective family $L_{V}\to V$ over an irreducible variety $V$ with $\dim V=\dim\mathrm{H}^{1}(L,\operatorname{\mathcal{O}}_{L})$ and a morphism $r_{V}\colon L_{V}\to M$ such that
  1. (1)
    there exists a point $0\in V$ such that $r_{0}=r\colon L=L_{0}\to M$,
  2. (2)
    $r_{s}\colon L_{s}\to M$ is birational onto its image and is an immersed Lagrangian for each $s\in V$, and
  3. (3)
    for any general point $s_{0}\in V$, the set of points $s\in V$ such that $r_{s}(L_{s})=r_{s_{0}}(L_{s_{0}})\subset M$ is finite.
Proof.
By the adjunction of functors, we have
$\operatorname{Ext}_{M}^{1}(r_{*}\mathcal{L},r_{*}\mathcal{L})=\operatorname{Ext}_{L}^{1}(r^{*}r_{*}\mathcal{L},\mathcal{L}).$
Since $\mathcal{H}^{i}(r^{*}r_{*}\mathcal{L})=0$ for any $i>0$, we have an exact sequence
$0\to\operatorname{Ext}^{1}_{L}(\mathcal{L},\mathcal{L})\to\operatorname{Ext}_{L}^{1}(r^{*}r_{*}\mathcal{L},\mathcal{L})\to\operatorname{Hom}_{L}(\mathcal{H}^{-1}(r^{*}r_{*}\mathcal{L}),\mathcal{L})\to\operatorname{Ext}^{2}_{L}(\mathcal{L},\mathcal{L}).$
This proves the first statement and gives an inclusion
$\operatorname{Ext}_{L}^{1}(r^{*}r_{*}\mathcal{L},\mathcal{L})\subset\operatorname{Ext}^{1}_{L}(\mathcal{L},\mathcal{L})\oplus\operatorname{Hom}_{L}(\mathcal{H}^{-1}(r^{*}r_{*}\mathcal{L}),\mathcal{L}).$ (15)
Since $\mathcal{L}$ is locally free and $L$ is normal, we get
$\operatorname{Hom}_{L}(\mathcal{H}^{-1}(r^{*}r_{*}\mathcal{L}),\mathcal{L})=\operatorname{Hom}_{L}((\mathcal{H}^{-1}(r^{*}r_{*}\mathcal{L}))^{\vee\vee},\mathcal{L}),$ (16)
where $(\mathcal{H}^{-1}(r^{*}r_{*}\mathcal{L}))^{\vee\vee}$ denotes the double dual (reflexive hull) of the sheaf $\mathcal{H}^{-1}(r^{*}r_{*}\mathcal{L})$.
Now, let $W\subset L$ be an open subset with $\mathrm{codim}_{L}(L\setminus W)\geq 2$ such that $r|_{W}$ is a closed embedding. Then by [36, Proposition 11.8], we have
$(\mathcal{H}^{-1}(r^{*}r_{*}\mathcal{L}))|_{W}\cong\mathcal{H}^{-1}((r|_{W})^{*}(r|_{W})_{*}\mathcal{L}|_{W})\cong(\mathcal{N}^{\vee}_{L/M}\otimes\mathcal{L})|_{W}.$
Therefore, using the fact that $\mathrm{codim}_{L}(L\setminus W)\geq 2$, the locally freeness of $\mathcal{N}^{\vee}_{L/M}\otimes\mathcal{L}$, and [76, Tag 0E9I], we have
$(\mathcal{H}^{-1}(r^{*}r_{*}\mathcal{L}))^{\vee\vee}\cong\mathcal{N}^{\vee}_{L/M}\otimes\mathcal{L},$
which together with $\mathcal{N}_{L/M}\cong\Omega_{L}$, (16), and (15) implies
$\operatorname{Ext}_{M}^{1}(r_{*}\mathcal{L},r_{*}\mathcal{L})=\operatorname{Ext}_{L}^{1}(r^{*}r_{*}\mathcal{L},\mathcal{L})\subset\mathrm{H}^{1}(L,\mathbb{C})$
and we get (14).
Finally, let $P\to V$ be the component of the relative Picard scheme of $L_{V}\to V$ containing the point corresponding to $\mathcal{L}$. Then we have
$\dim P=\dim V+\dim\mathrm{H}^{1}(L,\operatorname{\mathcal{O}}_{L})=\dim\mathrm{H}^{1}(L,\mathbb{C})$
by our assumption. Since $r_{s}$ is birational onto its image, we know that $r_{s_{*}}\mathcal{L}_{s}$ is a stable sheaf on $M$ for each $s\in V$ and line bundle $\mathcal{L}_{s}$ on $L_{s}$. Hence, up to shrinking $V$ if necessary, $r_{V}$ induces a morphism $P\to N$, where $N$ is an irreducible component of the moduli space of stable sheaves on $M$ such that $[r_{*}\mathcal{L}]\in N$. By the assumption (3), we know that $P\to N$ has finite general fibers, which gives $\dim N\geq\dim P=\dim\mathrm{H}^{1}(L,\mathbb{C})$. Therefore, we obtain
$\dim\operatorname{Ext}_{M}^{1}(r_{*}\mathcal{L},r_{*}\mathcal{L})=\dim T_{[r_{*}\mathcal{L}]}N\geq\dim N\geq\dim\mathrm{H}^{1}(L,\mathbb{C}),$
which implies the equality in (14). ∎
@@ 6.1. Double dual EPW sextics
Recall that by [18, Theorem 3.10], each smooth GM variety $X$ of dimension $\dim X\geq 3$ is uniquely determined by its Lagrangian data
@@
as in [61, Section 2]. See also [18, Appendix B] for an overview. We set $\mathsf{Y}_{A(X)}:=\mathsf{Y}_{A(X)}^{\geq 1}$ and $\mathsf{Y}_{A(X)^{\perp}}:=\mathsf{Y}_{A(X)^{\perp}}^{\geq 1}$, which are called the Eisenbud–Popescu–Walter (EPW) sextic and the dual EPW sextic associated with $X$, respectively.
Moreover, by [63, Section 1.2] and [20, Theorem 5.2], as $A(X)$ does not contain decomposable vectors [^8] This means $\mathbb{P}(A(X))\cap\operatorname{Gr}(3,V_{6})=\varnothing$, which holds for $A(X)$ when $X$ is smooth and $\dim X\geq 3$ by [18, Theorem 3.16]., we have natural double coverings
@@
In the rest of this section, we will mainly focus on (double) dual EPW varieties, as they exhibit a clearer connection to GM varieties.
@@
Since $X$ and $Y$ are both ordinary, by [18, Propoition 3.14], the Lagrangian data of $X$ and $Y$ are $(V_{6},V_{5},A(X))$ and $(V_{6},V_{5},A(Y))$ respectively, with $\dim(A(X)\cap A(Y))=9$. Therefore, part (1) follows directly from the above definition of EPW varieties.
@@
Consequently, $\iota$ is finite and unramified in codimension one. Thus, the natural map $W\to\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ is also unramified in codimension one. Then, by the smoothness of $\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ and [32, Theorem 1.6], it follows that $W\to\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ is unramified, and so is $\iota$.
@@
Moreover, since $W\to\mathsf{Y}_{A(Y)^{\perp}}^{\geq 2}$ is unramified and finite of degree two, by [76, Tag 0B3D] and [76, Tag 00R4], it is flat and hence an étale double covering. In particular, $W$ is smooth.
Finally, to show $W\cong\widetilde{\mathsf{Y}}_{A(Y)^{\perp}}^{\geq 2}$ and the diagram in (2) commutes, we use the uniqueness part of [20, Theorem 5.2(2)]. We first recall some notions in [20, Section 5.1]. Let $(V_{6},V_{5},A)$ and $(V_{6},V_{5},A^{\prime})$ be two Lagrangian data such that $\dim(A\cap A^{\prime})=9$ and both $A$ and $A^{\prime}$ are general. We define three bundles by $\mathscr{A}_{1}:=A\otimes\operatorname{\mathcal{O}}_{\mathbb{P}(V_{6})}$, $\mathscr{A}_{1}^{\prime}:=A^{\prime}\otimes\operatorname{\mathcal{O}}_{\mathbb{P}(V_{6})}$, and $\mathscr{A}_{2}:=\wedge^{2}T_{\mathbb{P}(V_{6})}(-3)$. These bundles are Lagrangian subbundles of the trivial symplectic bundle $\mathscr{V}:=\wedge^{3}V_{6}\otimes\operatorname{\mathcal{O}}_{\mathbb{P}(V_{6})}$ on $\mathbb{P}(V_{6})$. As in [20, Section 4], the symplectic structure induces injective maps
$\omega_{\mathscr{A}_{1},\mathscr{A}_{2}}\colon\mathscr{A}_{1}\to\mathscr{A}_{2}^{\vee}$
and
$\omega_{\mathscr{A}_{1}^{\prime},\mathscr{A}_{2}}\colon\mathscr{A}_{1}^{\prime}\to\mathscr{A}_{2}^{\vee}$
We set $\mathscr{C}:=\mathrm{cok}(\omega_{\mathscr{A}_{1},\mathscr{A}_{2}})$ and $\mathscr{C}^{\prime}:=\mathrm{cok}(\omega_{\mathscr{A}_{1}^{\prime},\mathscr{A}_{2}})$. Then as in [20, Section 5], $\mathsf{Y}^{\geq k}_{A}$ and $\mathsf{Y}^{\geq k}_{A^{\prime}}$ are exactly the rank $\geq k$ loci of sheaves $\mathscr{C}$ and $\mathscr{C}^{\prime}$, respectively.
Let $U:=\mathsf{Y}^{\geq 2}_{A^{\prime}}\setminus\mathsf{Y}^{\geq 2}_{A}$. We claim that $\mathscr{R}_{1}|_{U}\cong\mathscr{R}^{\prime}_{2}|_{U}$, where $\mathscr{R}_{1}:=(\mathscr{C}|_{\mathsf{Y}_{A}})^{\vee\vee}$ and $\mathscr{R}^{\prime}_{2}:=(\wedge^{2}(\mathscr{C}^{\prime}|_{\mathsf{Y}^{\geq 2}_{A^{\prime}}}))^{\vee\vee}$. Note that $\mathscr{C}|_{U}\cong\mathscr{R}_{1}|_{U}$ is a line bundle on $U$ by $U\subset\mathsf{Y}^{1}_{A}$ and $\mathscr{R}^{\prime}_{2}$ is also a line bundle since $\mathsf{Y}^{\geq 3}_{A^{\prime}}=\varnothing$. To prove the claim, we consider
$\mathscr{F}:=\mathrm{cok}(A\cap A^{\prime}\otimes\operatorname{\mathcal{O}}_{\mathbb{P}(V_{6})}\to\mathscr{A}_{2}^{\vee}),$
which fits into two exact sequences
$0\to\operatorname{\mathcal{O}}_{\mathbb{P}(V_{6})}\to\mathscr{F}\to\mathscr{C}\to 0$ (17)
and
$0\to\operatorname{\mathcal{O}}_{\mathbb{P}(V_{6})}\to\mathscr{F}\to\mathscr{C}^{\prime}\to 0$ (18)
From the definition, we know that the rank $\geq 3$ locus of $\mathscr{F}$ can be described as
$\{[v]\in\mathbb{P}(V_{6})\mid\mathrm{dim}(A\cap A^{\prime}\cap(v\wedge\bigwedge^{2}V_{6}))\geq 2\},$
which is contained in $\mathsf{Y}^{\geq 2}_{A^{\prime}}\cap\mathsf{Y}^{\geq 2}_{A}$. Therefore, the fiber of $\mathscr{F}$ has rank $\leq 2$ at each point of $U$ by the definition of $U$. As $\mathscr{C}^{\prime}|_{\mathsf{Y}^{\geq 2}_{A^{\prime}}}$ is locally free of rank $2$ by definition, then by restricting the exact sequence (18) to $U$, we get an isomorphism $\mathscr{F}|_{U}\cong\mathscr{C}^{\prime}|_{U}$. In particular, $\mathscr{F}|_{U}$ is a bundle of rank $2$ on $U$. On the other hand, as $\mathsf{Y}^{\geq 2}_{A^{\prime}}\neq\mathsf{Y}^{\geq 2}_{A}$, we know that $\mathscr{C}|_{\mathsf{Y}^{\geq 2}_{A^{\prime}}}$ has generic rank $1$ on $\mathsf{Y}^{\geq 2}_{A^{\prime}}$ as $\mathsf{Y}^{\geq 2}_{A^{\prime}}$ is integral by [20, Theorem 5.1]. Therefore, the kernel $\mathscr{F}|_{U}\twoheadrightarrow\mathscr{C}|_{U}$ has generic rank $1$ on $U$ and the pull-back of (17) gives an exact sequence
$0\to\operatorname{\mathcal{O}}_{U}\to\mathscr{F}|_{U}\to\mathscr{C}|_{U}\to 0$
Combing this with $\mathscr{F}|_{U}\cong\mathscr{C}^{\prime}|_{U}$ implies
$\mathscr{R}^{\prime}_{2}|_{U}=\wedge^{2}(\mathscr{C}^{\prime}|_{U})\cong\mathscr{C}|_{U}=\mathscr{R}_{1}|_{U}.$
This proves the claim.
Now, back to the proof of (2). We set $A:=A(X)^{\perp}$ and $A^{\prime}:=A(Y)^{\perp}$ as above. By the uniqueness part of [20, Theorem 5.2(2)], we need to show the étale covering $\gamma\colon W\to\mathsf{Y}_{A^{\prime}}^{\geq 2}$ satisfies
$\mathrm{cok}(\operatorname{\mathcal{O}}_{\mathsf{Y}_{A^{\prime}}^{\geq 2}}\to\gamma_{*}\operatorname{\mathcal{O}}_{W})\cong\mathscr{R}^{\prime}_{2}(-3),$
where $\mathscr{R}^{\prime}_{2}:=(\wedge^{2}(\mathscr{C}^{\prime}|_{\mathsf{Y}^{\geq 2}_{A^{\prime}}}))^{\vee\vee}$. Since $\gamma$ coincides with the base change of
$\widetilde{\mathsf{Y}}_{A}\to\mathsf{Y}_{A}$
along $\mathsf{Y}^{\geq 2}_{A^{\prime}}\hookrightarrow\mathsf{Y}_{A}$ outside $\mathsf{Y}^{\geq 2}_{A}$, by [20, Theorem 5.2(1)] and the base change theorem, we have
$(\mathscr{R}_{1}(-3))|_{U}=((\mathscr{C}|_{\mathsf{Y}_{A}}(-3))^{\vee\vee})|_{U}\cong\mathrm{cok}(\operatorname{\mathcal{O}}_{\mathsf{Y}_{A^{\prime}}^{\geq 2}}\to\gamma_{*}\operatorname{\mathcal{O}}_{W})|_{U}$
for $U:=\mathsf{Y}^{\geq 2}_{A^{\prime}}\setminus\mathsf{Y}^{\geq 2}_{A}$. Then, from the claim we just proved, this implies
$\mathscr{R}^{\prime}_{2}(-3)|_{U}\cong\mathrm{cok}(\operatorname{\mathcal{O}}_{\mathsf{Y}_{A^{\prime}}^{\geq 2}}\to\gamma_{*}\operatorname{\mathcal{O}}_{W})|_{U}.$
From $\dim(\mathsf{Y}_{A^{\prime}}^{\geq 2}\setminus U)=0$ by Lemma 6.1, the normality of $\mathsf{Y}_{A^{\prime}}^{\geq 2}$, and [76, Tag 0AY6], we obtain
$\mathscr{R}^{\prime}_{2}(-3)\cong\mathrm{cok}(\operatorname{\mathcal{O}}_{\mathsf{Y}_{A^{\prime}}^{\geq 2}}\to\gamma_{*}\operatorname{\mathcal{O}}_{W}).$
This finishes the proof of (2). ∎
@@
It is suggested in [40, Section 5.1] that $i$ is injective, implying that $i$ is an embedding. Nevertheless, this is equivalent to $\varnothing=\mathsf{Y}^{\geq 2}_{A(Y)^{\perp}}\cap\mathsf{Y}_{A(X)^{\perp}}^{\geq 2}\subset\mathsf{Y}_{A(X)^{\perp}}$, which is impossible as explained in [25, Lemma 3.2]. See Remark 6.15 for an alternative proof of the non-injectivity of $i$.